find n factor (NH 4 ) 2 Cr­ 2 O 7 -----> N 2 + Cr 2 O 3 +4H 2 O

find n factor


(NH4)2Cr­2O7----->N2 + Cr2O3 +4H2O

Grade:12th pass

1 Answers

sumukh purohit
13 years ago

Ox No. of N=-3 therefore change in ox. no=-3-0=-3

n factor = 2*3=6